OpenStudy (anonymous):

A 25000 kg bowling ball moving at 40 m/s this a 10000 kg bowling ball moving at 25 m/s from behind. Assuming a one-dimensional perfectly elastic collision, what are the final velocities of the two balls?

OpenStudy (anonymous):

Use energy and momentum conservation. p1(initial)+p2(initial)=p1(final)+p2(final) E(initial)=E(final) => p1(i)²/2*m1+p2(i)²/2*m2=p1(f)²/2*m1+p2(f)²/2*m2 You have the values of initial momenta and kinetic energy. So you have a system of two equation and two unknows (p1(f) and p2(f))